OpenStudy (anonymous):

find dy/dx such that dx/dy =dy/dx if \[y=x^3\]

OpenStudy (anonymous):

I meant find x.

OpenStudy (turingtest):

dy/dx=3x^2 x=y^(1/3) dx/dy=(1/3)y^(-2/3) dx/dy=dy/dx (1/3)y^(-2/3)=3x^2 x^2=(1/9)y^(-2/3) x=(1/3)y^(-1/3)